Determine the concavity of the parabola representing the quadratic function, its y intercept, its x
intercept if any exist, and the coordinates of the vertex. Sketch the parabola. Y = 25x
2 +8x-12.
1
Expert's answer
2020-05-08T15:12:49-0400

Let us consider a parabola

"y(x) =ax^2+bx+c."

If "a>0", then for very large "x" we get "y(x) >0," analogous for negative "x"with large modulus. Therefore, if "a>0",then parabola is open (concave) up. In our example "a=25," so the parabola is concave up.


Now let us determine the vertex. The x-coordinate of vertex is

"x_0=\dfrac{-b}{2a}=\dfrac{-8}{2\cdot 25} = - 0.16."

The second coordinate can be calculated as

"y_0=y(x_0)=25\cdot(-0.16)^2+8\cdot (-0.16)-12=0.64-1.28-12=-12.64."


Next we obtain the intersection point of the parabola and y-axis. It is the point "(0, y(0))" or "(0, - 12)."


Next we'll determine the intersection points of the parabola and x-axis. We shoul solve an equation

"25x^2+8x-12=0."

"D=8^2+4\cdot25\cdot12 = 1264 =2^4\cdot79."

Therefore, roots of equation are

"x_1 = - \dfrac{4+2\sqrt{79}} {25} ,\; x_2 = - \dfrac{4-2\sqrt{79}} {25}."